The period T(time in seconds for one complete cycle) of a simple pendulum is related to the length L (in feet) of the pendulum by the formula 8T^2=pi^2L. If a child is on a swing with a 10-foot chain, then how long does it take to complete one cycle of the swing?
Answer by checkley77(12844)

8T^2=pi^2L
8t^2=9.86*10
8t^2=98.6
t^2=98.6/8
t^2=12.345
t=sqrt12.345
t=3.51 seconds.